OpenStudy (anonymous):

Two equations are shown: • Equation 1: 3 4 (x−12)=12 •Equation 2: 3 4 y−12=12

OpenStudy (anonymous):

x=17.625?

OpenStudy (anonymous):

Solving for x and y?

OpenStudy (anonymous):

Your question isn't very clear

OpenStudy (anonymous):

solve for x and y. provide the equation be true

OpenStudy (anonymous):

Two equations are shown: • Equation 1: 3 4 (x−12)=12 •Equation 2: 3 4 y−12=12 Solve each equation. Then, enter a number in each box to make this statement true. The value of x is: The value of y is:

OpenStudy (anonymous):

What are 3 and 4, still not clear, just use draw or write it out yourself, copying and pasting the same thing isn't helpful at all..

OpenStudy (anonymous):

3/4(x-12)=12

OpenStudy (anonymous):

3/4y-12=12

OpenStudy (anonymous):

@iambatman i keep getting different answer each time i work it

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }(x-12)=12 \implies \frac{ 3 }{ 4 }x-9=12\] can you solve x from here, you have to distribute the 3/4 over x and -12

OpenStudy (anonymous):

i got that part with = 8.25

OpenStudy (anonymous):

8.25? How?

OpenStudy (anonymous):

12+9 then divide 3/4

OpenStudy (anonymous):

No you multiply both sides by 4/3

OpenStudy (anonymous):